OpenStudy (anonymous):

The width of a rectangle is 5 less than twice its length. If the area of the rectangle is 85 cm, what is the length of the diagonal?

OpenStudy (campbell_st):

well let the length = l then the width is 2l - 5 then 85 = l (2l - 5) distributing gives 85 = 2l^2 - 5l or the quadratic \[2l^2 - 5l - 85 = 0\] solve the quadratic for l

OpenStudy (anonymous):

I cant seem to find the answer?

OpenStudy (anonymous):

\[Area=85cm^2\] \[Width=(2*length)-5\] Now the area is determined by \[Area=length*width\] Which gives you: \[85=length*(2length-5)=2length^2-5length \rightarrow 0=2length^2-5length-85\] Solve the equation and find the length. When you have done this, just use the pythagorean formula to find the length of the diagonal. Since the diagonal is determined by \[diag=\sqrt{length^2+width^2}\] You'll just have to put your numbers in and you will get your answer.